# Expert Router: Orchestrating Efficient Language Model Inference through Prompt Classification

### v1

@@ IV Experimental Setup
This section introduces the experimental setup for inference testing, including a description of the models employed within the Expert Router and the chosen baseline model. It also describes the approach for simulating high-load scenarios by increasing the number of concurrent users, realized through the User Simulator in Figure 1. These experiments aim to assess the Expert Router’s performance and benchmark it against a high-throughput baseline model.
IV-A Model deployment
This study utilizes LLama 2 models [3] across all experiments. All experiments are conducted on a DGX H100 system with eight H100 GPUs. Each of them features 80 GB of GPU memory. The experimental framework is given by two primary scenarios: one involving runs utilizing a baseline model and the other our Expert Router architecture.
The aim is to evaluate the Expert Router’s performance by benchmarking it against a baseline model optimized for high throughput. This baseline model is a Llama 2 configuration with 70 billion parameters (70B), tensor parallelized over eight GPUs, and FP16 weights.
A) 70B FP16 TP8
As outlined in Table I, the baseline model operates with a large batch size of 600, optimized to maximize throughput. For the baseline results, the user requests are directly sent to the Triton inference server hosting the tensor-parallel 70B model without the routing gateway and load balancer as intermediate layers.
@@ IV-A Model deployment
TABLE I: Key specifications of the Models For all experiments conducted within this study, LLama 2 models are used. Besides the batch size, the table lists memory-related specifications for each of the models, such as the weight and the KV-cache size in GB.
@@
Baseline Model
A) 70B FP16 TP8 141 GB 366 GB 600
Orchestrated Models
B) 70B INT8 (Exp. Router) 69 GB 12 GB 60
C) 13B FP16 (Exp. Router) 26 GB 54 GB 40
D) 13B FP8 (Exp. Router) 13 GB 65 GB 150
Our research examines the impact of varying model specifications and configurations on system performance for the models integrated within the Expert Router. Specifically, we investigate three configurations: the first (B) uses eight quantized 70B models to explore high-parameter scenarios; the second (C) involves eight 13 billion parameters (13B) models, allowing for a larger KV-cache and FP16 weights; and the third (D) employs eight 13B models optimized for the H100 hardware using FP8 format. With every specification, our objective is to optimize the batch size within the constraints of available GPU memory, ensuring maximal utilization.
B) 70B INT8 (Exp. Router)
@@
The first configuration uses a Llama 2 70B model with INT8 quantized weights and KV-cache. We employ weight-only quantization provided within the TensorRT-LLM library [46] to build the model. Table I shows that the resulting model uses 69 GB of GPU memory for its weights and $\mathrm{72}\times\mathrm{10^{3}}$ tokens in the KV-cache, allocating 12 GB of GPU memory. The batch size is set to 60. Each GPU hosts a single INT8 70B model in this experiment, leading to 560 billion INT8 quantized weights across the system. In our tests, we included this particular model setup to explore how the Expert Router performs when handling models with many parameters.
C) 13B FP16 (Exp. Router)
@@
The second configuration uses a Llama 2 13B model, requiring 26 GB for weight storage. This reduction in memory requirement for weights allows allocating a greater portion of available GPU memory to the KV-cache. In this instance, the KV-cache accommodates $\mathrm{160}\times\mathrm{10^{3}}$ tokens. This allows for the evaluation of the Expert Router’s performance with non-quantized models that leverage a larger KV cache.
D) 13B FP8 (Exp. Router)
Lastly, we deploy a quantized version of the 13B model, with both weights and the KV-cache formatted in FP8. We utilized the NVIDIA Algorithmic Model Optimization (AMMO) toolkit for weight quantization, part of TensorRT-LLM [46]. This approach allows us to explore the impact of an expanded KV-cache on system performance and leverage the computational efficiency offered by the NVIDIA H100’s enhanced support for FP8 matrix multiplications [52]. In this setup, each model accommodates approximately $\mathrm{159}\times\mathrm{10^{3}}$ tokens in its KV-cache, testing the Expert Router´s capability to manage models with both a large KV-cache and an efficient data type.
To further improve inference performance, each model incorporates the same optimization techniques. Among these, XQA kernels are employed, which improve Multi-query Attention (MQA) and Grouped-query Attention (GQA) processes during the generation phase [46]. MQA is a modification of the multi-head attention mechanism used in Transformers, where all attention heads share the same set of Key and Value matrices. This design significantly reduces memory bandwidth requirements during incremental inference [53]. GQA modifies the attention mechanism by organizing multiple attention heads into groups, each sharing resources for analyzing segments of the incoming sequences. This structure decreases the number of operations required for processing these sequences [54]. Additionally, all models employ a paged attention mechanism, which enhances the memory efficiency [55].
IV-B User Simulator
The user simulator manages N synthetic users, sending prompts concurrently to emulate the increasing load on the inference systems. This is accomplished by spawning N independent threads, each functioning independently as a user. These users draw their prompts from a selected segment of the C4 dataset, with a uniform probability distribution across eight categories defined by Gurungan et al. [49] [18]. This ensures an equal distribution of computational workload across the system’s clusters.
The rationale behind the uniform distribution is that it aligns with the scalability of the Expert Router. Such systems can dynamically and autonomously adjust by adding more models to balance spikes in demand for individual clusters, thereby maintaining uniformity in request distribution [56]. The C4 dataset was selected due to its use by Gurungan et al. in developing their k-means clustering algorithm [18]. Each synthetic user generates a prompt, transmitted to the backend via gRPC, with a token limit of 300 $\pm$ 30. Moreover, each user requests a response prompt of 1200 tokens, a condition enforced by continuous token requests even after the model has indicated sequence completion.
Inference operations are conducted in streaming mode, allowing for measuring the arrival times of individual tokens. This data is stored and later used in the performance evaluation presented in Section V. To simulate a chat scenario, each user sends two sequential prompts to the backend, including the history of the preceding interaction. The initial prompt will be referred to as the ”base prompt,” while the subsequent prompt will be termed the ”continuation prompt.” The base prompt contains around 335 $\pm$ 30 input tokens, whereas the continuation prompt includes about 1535 $\pm$ 30 tokens. This allows to investigate the system under two different load scenarios.
In this study, we increase the number of users from a single user to 1,000 in increments of 100. Furthermore, each run is repeated ten times to measure the uncertainty of the derived metrics.
IV-C Infrastructure Setup
$10^{0}$
$10^{1}$
$10^{2}$
$10^{3}$
Time to First Token (s)
Base Prompt
1
100
200
300
400
500
600
700
800
900
1000
Refer to captionRefer to captionRefer to caption
A) 70B FP16 TP8
B) 70B INT8 (Exp. Router)
C) 13B FP16 (Exp. Router)
D) 13B FP8 (Exp. Router)
Nr. Users
Figure 2: Time to First Token for the base prompt: The models orchestrated via the Expert Router (B-D) exhibit a significantly larger TTFT compared to the tensor parallelized baseline model (A). This is caused by the beneficial parallelized computing of the input sequence by the baseline model (A) during the prefill phase.
1
1
100
200
300
400
500
600
700
800
900
1000
Nr. Users
$10^{0}$
$10^{1}$
$10^{2}$
$10^{3}$
Time to First Token (s)
Continuation Prompt
A) 70B FP16 TP8
B) 70B INT8 (Exp. Router)
C) 13B FP16 (Exp. Router)
D) 13B FP8 (Exp. Router)
Figure 3: Time to First Token for continuation prompt: Similar to the base prompt, the orchestrated models (B-D) exhibit a significantly higher TTFT compared to the tensor parallelized model (A). No drastic increase of the TTFT between the base and continuation prompt is visible even though the input sequence length has increased.
For all of our experiments, we use a DGX H100. We deploy the Llama 2 70B model in the baseline configuration (A in Table I) through a Triton server, employing tensor parallelization across all eight GPUs. This method is intended to maximize the computational resources available on the DGX system for high-throughput inference tasks.
For the Expert Router experiments, the architecture hosts both the k-means clustering algorithm and individual Triton containers on the same GPUs. Specifically, we deploy 16 instances of the routing gateway, each requiring 1,138 MB of GPU memory, resulting in 2,276 MB on each GPU dedicated to the Expert Router system. The remaining GPU memory is allocated for deploying the individual expert models.
Additionally, the user simulator component is deployed on the CPUs of the same DGX H100 system. By centralizing the experimental setup on a single DGX H100, we aim to reduce potential distortions caused by network latency. However, this approach might lead to interferences among the different components, including the Expert Router, the Triton servers, and the User Simulator, due to the shared use of the system’s resources. Such interferences are expected to have a minimal impact on the overall performance metrics.
@@ V Evaluation Methodology
Throughout our experiments, we assess the performance of the Expert Router and the baseline model from two perspectives: user- and system-centric Metrics. This strategy allows us to evaluate system-relevant metrics and analyze the effects of different model configurations on the user interaction latency and throughput.
@@ V-A User-Centric Metrics
In the evaluation from the user perspective, we present three metrics. The first is the Time to First Token (TTFT) [19]. It measures the duration from when the user sends a request to when the user receives the first token. We report the TTFT average with increasing number of users. Adjacent to the TTFT, we track the Time Per Output Token (TPOT). It describes the time difference between consecutive tokens. Again, we report the average TPOT for an increasing number of users. Finally, we also investigate the throughput from the user perspective. The user throughput measures the rate at which tokens are received, as experienced by a single user. This metric is calculated by dividing the total number of response tokens by the duration from the request’s start ($t_{s}$) to the completion of the answer ($t_{e}$). We average this value across all users and report the mean and standard deviation.
@@ V-B System-Centric Metrics
Base Prompt
$10^{0}$
Time per Output Token (s)
1
100
200
300
400
500
600
700
800
900
1000
Nr. Users
$10^{-1}$
$10^{-2}$
A) 70B FP16 TP8
B) 70B INT8 (Exp. Router)
C) 13B FP16 (Exp. Router)
D) 13B FP8 (Exp. Router)
Figure 4: Time Per Output Token for the base prompt: For concurrent user counts above 200, the tensor parallelized baseline model (A) exhibits a higher median TPOT than the orchestrated models (B-D). This can be attributed to the communication cost between the eight GPUs.
Continuation Prompt
1
100
200
300
400
500
600
700
800
900
1000
Time per Output Token (s)
$10^{0}$
$10^{-1}$
Nr. Users
A) 70B FP16 TP8
B) 70B INT8 (Exp. Router)
C) 13B FP16 (Exp. Router)
D) 13B FP8 (Exp. Router)
Figure 5: Time Per Output Token for the continuation prompt: Similar to the base prompt, for concurrent user counts above 200, the tensor-parallelized baseline model (A) exhibits a higher TPOT compared to the orchestrated models (B-D).
Additionally, we adopt two methods to calculate the system’s total throughput: session throughput and time-bucket throughput.
Session Throughput
The first approach measures the time it takes for all users to receive their responses in a single session. Again, we use the p99 response time to measure the upper latency bounds under the specific load scenario. The number of users receiving responses within this p99 interval is $n_{99}$. For this group, we define the session’s start and end times by the earliest and latest response times recorded among these users.
$t_{S}$ $=\mathrm{min}(t_{s,1},t_{s,2},...,t_{s,n_{99}})$ (1)
$t_{E}$ $=\mathrm{max}(t_{e,1},t_{e,2},...,t_{e,n_{99}})$ (2)
The throughput for a specific number of concurrent users can be calculated by following equation
$Th_{n_{99}}=\frac{1}{(t_{E}-t_{S})}\sum_{u=1}^{n_{99}}m_{u}$ (3)
where $m_{u}$ is the total number of tokens user $u$ received in the answer prompt. The throughput value obtained from Equation 3 enables direct comparison between the different systems deployed in this study. However, this value provides only a broad average, not offering detailed insights into the system’s throughput behavior. For a more precise analysis, we track the throughput changes over time.
Time Bucket Throughput
To add another perspective to our analysis, we assess each trial based on the timestamps associated with all tokens. In each trial, we establish K time intervals (or buckets) and count the number of tokens present within each interval. Thus, we only count how many tokens the system produces in a specific time interval, regardless to which user a specific token belongs. Significantly, this approach does not depend upon the start and end times for the throughput analysis. Equation 3 illustrates that the derived throughput can be significantly influenced by the span between the start and end times. By adopting the time bucket strategy, we mitigate this sensitivity, focusing solely on the distribution of token timestamps.
@@ VI Results
0
200
400
600
800
1000
Nr. Users
Throughput (token/s)
0
200
400
600
800
1000
Nr. Users
Base Prompt
Continuation Prompt
$10^{0}$
$10^{1}$
$10^{2}$
A) 70B FP16 TP8
B) 70B INT8 (Exp. Router)
C) 13B FP16 (Exp. Router)
D) 13B FP8 (Exp. Router)
Refer to captionRefer to caption
$10^{0}$
$10^{1}$
$10^{2}$
Figure 6: User Throughput: For more than one concurrent user, no clear difference between the models orchestrated by the Expert Router (B-D) and the tensor-parallelized baseline model (A) is visible.
0
200
400
600
800
1000
Nr. Users
0
200
400
600
800
1000
1200
1400
p99 Response time (s)
A) 70B TP8
B) 70B INT8 (Exp. Rouer)
C) 13B FP16 (Exp. Router)
D) 13B FP8 (Exp. Router)
0
200
400
600
800
1000
Nr. Users
0
200
400
600
800
1000
1200
1400
Base Prompt
Continuation Prompt
Figure 7: p99 Response time: The p99 response time is smaller for Expert Router orchestrated models (B-D) across user counts of 200-900 compared to the tensor-parallelized baseline model (A).
In this section, we present the outcomes of our experiments, analyzing the performance from both user and system perspectives. To assess the performance of our proposed inference architecture under varied load conditions, we systematically increased the number of concurrent users sending requests to the system. Additionally, we compare the performance between the base prompt and the continuation prompt.
In addition to the performance metrics presented in Section V we have also measured the latency introduced by the routing gateway. The average latency for each user when sending prompts via the routing gateway is 442 ms with a standard deviation of 302 ms. This high standard deviation is caused by outliers with very high latencies, which could originate from a bottleneck on the DGX system during peak load times. However, as the overall response times for inference are much larger, the latencies introduced by the routing gateway are negligible.
@@ VI-A User-Centric Metrics
Base Prompt
Continuation Prompt
0
200
400
600
800
1000
1200
1400
p99 Response Time (s)
0
200
400
600
800
1000
1200
Throughput (token/s)
1
400
700
200
500
600
800
900
1000
300
1000
100
300
400
500
600
700
800
900
A) 70B FP16 TP8
B) 70B INT8 (Exp. Router)
C) 13B FP16 (Exp. Router)
D) 13B FP8 (Exp. Router)
0
200
400
600
800
1000
1200
1400
p99 Response Time (s)
0
200
400
600
800
1000
1200
1
300
500
700
200
400
600
800
900
1000
1000
100
300
400
500
600
700
800
900
Figure 8: Session throughput: The session throughput, as computed by Equation 3, is higher for Expert Router orchestrated models (B-D) across user counts of 200-900 compared to the tensor-parallelized baseline model (A).
Window Number
Throughput (token/s)
0
10
20
30
$10^{0}$
$10^{1}$
$10^{2}$
$10^{3}$
$10^{4}$
1 User
100 User
300 User
600 User
1000 User
0
10
20
30
0
10
20
30
Window Number
A) 70B FP16 TP8
B) 70B INT8 (Exp. Router)
Window Number
C) 13B FP16 (Exp. Router)
D) 13B FP8 (Exp. Router)
0
10
20
30
40
Window Number
Figure 9: Time bucket throughput for the base prompt. The graphs display the time bucket throughput per model, with time windows of 60s. Each line representing a specific user concurrency level. The tensor-parallelized baseline model (A) exhibits sharp peaks, whereas the models orchestrated by the Expert Router (B-D) demonstrate consistent fluctuations in throughput resulting in a higher average throughput.
Figure 2 and 3 illustrates the Time-to-First-Token (TTFT) for the base and continuation prompt. The results demonstrate a notable performance gap between the tensor-parallelized baseline model (A) and those orchestrated by the Expert Router (B-D). The tensor-parallelized model (A) significantly outperforms the systems orchestrated via the Expert Router. This discrepancy can be attributed to the tensor-parallelized model’s ability to distribute the first stage of inference computations across multiple GPUs. In this stage, the KV-cache is populated by the input’s K and V vectors. This prefill phase involves processing input tokens through a highly parallelizable matrix-matrix multiplication [41]. Each GPU processes a segment of this computation independently and in parallel, and these results are subsequently aggregated through an all-reduce operation for generating the first output token [57]. Due to this parallel execution of the prefill phase, the tensor-parallelized model (A) achieves a significantly shorter TTFT.
The similarity between the base and continuation prompt indicates that the increase in load, due to the increased number of input tokens, does not significantly affect the performance. However, the Expert Router coordinated models (B-D) show decreased variance in the continuation prompt due to this consistent load across all users.
The TTFT advantage of the tensor-parallelized baseline model (A) turns into a disadvantage for the Time Per Output Token (TPOT) as shown in Figures 4 and 5. Specifically, the TPOT increases for the tensor-parallelized model (A) from 0.1 seconds at 200 users to nearly 1 second at 1000 users for both prompt types. In contrast, the orchestrated models (B-D) maintain a stable TPOT of about 0.1 seconds across user counts. Again, no clear distinction between the different orchestrated models can be seen. By working independently on each GPU, the orchestrated models eliminate the need for inter-GPU communication, thus embracing an embarrassingly parallel framework. This results in a significantly reduced inter-token latency.
Figure 6 evaluates the user throughput in response to increasing system load. This throughput depends on the total response time which can be calculated by $\mathrm{TTFT}+\mathrm{TPOT}\times\mathrm{Nr}_{\mathrm{tokens}}$. As the number of received tokens is constant, the behavior of the throughput observed by each user only depends on the TTFT and TPOT. Although the tensor-parallelized model (A) demonstrates a notably lower TTFT as the number of users increases, its latency between tokens increases. This increase in latency results in a throughput that, from the user’s perspective, closely aligns with that of the models orchestrated by the Expert Router (B-D). Conclusively, the scaling of the Expert Router is beneficial in scenarios where many response tokens are required as it increases the total response time less drastically.

### v2

@@ IV Experimental Setup
This section describes the experimental setup designed to evaluate the impact of the Expert Router on inference performance. We describe the characteristics of the models integrated into the Expert Router and the reason behind their selection. Additionally, we present the baseline models identified through preliminary experiments aimed at optimizing performance. The section also details the User Simulator to generate high-load scenarios, enabling a thorough comparison between the Expert Router and the baseline models.
IV-A User Simulator
The User Simulator manages N synthetic users, concurrently sending prompts to emulate increasing load on the inference systems. This is achieved by spawning N independent threads, each acting as a separate user. The simulated users draw their prompts from a selected segment of the C4 dataset [40], which are distributed across eight categories defined by Gurungan et al. [15]. We conduct experiments using two different request scenarios to analyze the Expert Router configurations.
The first one is a uniform distribution across the eight categories, such that each model needs to process the same amount of queries. The rationale behind the uniform distribution is that it aligns with the scalability of the Expert Router. Such systems can dynamically and autonomously adjust by adding more models to balance spikes in demand for individual clusters, thereby maintaining uniformity in request distribution [42].
In the second scenario, we partition the requests across the eight categories using a normal probability distribution. This results in some models receiving more inference requests than others, leading to an unbalanced load that increases the dynamics within the system.
Figure 2: Distribution of number of input tokens. The plot shows the distribution of the number of input tokens in the test set. Its shape is based on the input length distribution used by MLPerf in the Llama 2 70B benchmark test [26].
Figure 3: Preliminary Experiments: The two plots show the trajectories of the response times with increasing user concurrency for different batch sizes and data types. The results on the left side use a LLama 3 70B model parallelized over 4 GPUs (TP=4) and on the right side across 8 GPUs (TP=8). Based on these results we select the three baseline models listed in Table I.
The C4 dataset was selected due to its use by Gurungan et al. in developing their k-means clustering algorithm [15]. Each synthetic user generates a prompt, transmitted to the backend via gRPC. The distribution of input token lengths in our test dataset mirrors that of the dataset used by MLPerf in their Llama 2 70B inference benchmark test [26]. The token-length distribution across the test set is shown in Figure 2. This variability in input tokens introduces additional stress on the system, allowing us to investigate dynamic user behaviors more thoroughly.
Furthermore, we do not enforce a specific number of output tokens produced by the inference systems. The token generation is terminated with the arrival of the end of sequence token. However, we limit the maximum output length to 1000 tokens. Inference is carried out in streaming mode, allowing for measuring the arrival times of individual tokens. This data is stored and later used in the performance evaluation presented in Section V.
IV-B Infrastructure Setup
We use a DGX H100 for all experiments to host the baseline and expert router models. It contains eight H100 GPUs, each with 80 GB of GPU memory. To prevent interference between load generation and model inference, the User Simulator is deployed on a separate node. This node has an Intel Xeon Platinum 8480CL and two terabytes of RAM. The latency between the two nodes is measured at 17 $\pm$ 5 $\mu s$ and is negligible in the inference experiments.
We deploy the Llama 3 70B model in the baseline configurations through a Triton server, employing tensor-parallelization across eight and four GPUs. The specific deployment details will be described in the next section. For the Expert Router experiments, the architecture hosts the routing gateway and individual Triton containers on the same GPUs. Specifically, we deploy 16 instances of the routing gateway, each requiring 1,138 MB of GPU memory, resulting in 2,276 MB on each GPU dedicated to the Expert Router system. The remaining GPU memory is allocated for deploying the individual expert models.
IV-C Baseline Configurations
To accurately benchmark and compare the performance of the Expert Router, we must establish well-optimized baseline models. Therefore, we conduct preliminary experiments to determine the best configuration by testing various batch sizes ($B_{S}=\{20,100,200,400,600\}$), data formats (FP16, FP8, INT4), and levels of tensor-parallelism (TP=4, TP=8). We evaluate each configuration by increasing the number of concurrent users from 1 to 500 in steps of 100.
Given the extensive combination of hyperparameters that are being tested, we limit the number of answer tokens to 200 to reduce the runtime of each experiment. Additionally, we use a test set with 335 $\pm$ 30 input tokens. Each test run is repeated five times to estimate variances in response times. All preliminary experiments are conducted using the Llama 3 70B model.
For TP=4 configurations, deploying two instances of the model is possible on a DGX H100. Therefore, each model runs on four dedicated GPUs. The preliminary experiments measure response times based on a single model parallelized across 4 GPUs. Two instances will be deployed for the final results presented in Section VI, with requests distributed among the two models.
Fig 3 illustrates the response time trajectories as the number of users increases. The left plot shows the best-performing batch sizes for each data type under TP=4, while the right plot shows the results for TP=8.
For TP=8, the FP8 and FP16 models with a batch size 200 demonstrate similar performance when handling 500 concurrent users. However, for smaller numbers of users, the FP8 model outperforms the FP16 model. To ensure a comprehensive evaluation, both models will be included in the final experiments. We will refer to these two baseline models as A) 70B FP16 TP8 and B) 70B FP8 TP8 in the following sections. As shown in Table I, the reduced precision of baseline model B allows it to utilize over 100 GB more GPU memory for storing user context compared to baseline model A.
In the TP=4 results, the FP8 models consistently show a slight advantage. Therefore, in the final tests, we will deploy and evaluate two instances of the FP8 model with a batch size of 100 corresponding to baseline model C in Table I.
IV-D Expert Router Configuration
For the model configuration in the Expert Router, we utilize two different versions of the Llama3 model group: Llama3 70B and 8B. This approach allows us to examine how varying model specifications and configurations impact inference performance when integrated within the Expert Router. The first configuration (D) uses eight quantized 70B models to explore high-parameter scenarios; the second (E) involves eight billion parameters (8B) models, allowing for a larger KV-cache and FP16 weights. With every specification, our objective is to optimize the batch size within the constraints of available GPU memory, ensuring maximal utilization.
@@ IV-D Expert Router Configuration
TABLE I: Model Specifications: For all experiments conducted within this study, LLama 3 models are used. Besides the batch size, the table lists memory-related specifications for each of the models, such as the weight and the KV-cache size in GB.
@@
Baseline Models
A) 70B FP16 TP8 148.8 GB 347.28 GB 200
B) 70B FP8 TP8 69.6 GB 487.2 GB 200
C) 70B FP8 TP4 69.6 GB 221.2 GB 100
Expert Router Models
D) 70B INT4 (Exp.R.) 37.7 GB 37.6 GB 125
E) 8B FP16 (Exp.R.) 15.4 GB 62.2 GB 125
D) 70B INT8 (Exp.R.)
@@
The first configuration uses a Llama 3 70B model with INT4 quantized weights. We employ activation aware quantization (AWQ) [43] provided within the TensorRT-LLM library [37] to build the model. Table I shows that the resulting model uses 37.7 GB of GPU memory for its weights and allocates 37.6 GB of GPU memory for the KV-cache. The batch size is set to 125. Each GPU hosts a single INT4 70B model in this experiment, leading to 560 billion INT4 quantized weights across the system. In our tests, we included this particular model setup to explore how the Expert Router performs when handling models with many parameters.
E) 8B FP16 (Exp.R.)
@@
The second configuration uses Llama 3 8B model, requiring 15.4 GB for weight storage. This reduction in memory requirement for weights enables allocating a greater portion of available GPU memory to the KV-cache with 62.2 GB. This allows for evaluating Expert Router’s performance with non-quantized models that leverage a larger KV cache.
@@ V Evaluation Methodology
Throughout our experiments, we assess the performance of the Expert Router and baseline models from two perspectives: user-centric and system-centric metrics. This strategy allows us to evaluate system-relevant metrics and analyze the effects of different Expert Router configurations on the interaction latency and throughput.
@@ V-A User-Centric Metrics
In the evaluation from the user perspective, we present three metrics. The first is the Time to First Token (TTFT) [44]. It measures the duration from when the user sends a request to when the user receives the first token. We report the TTFT average with an increasing number of users. Adjacent to the TTFT, we track the Time Per Output Token (TPOT). It describes the time difference between consecutive tokens. Again, we report the average TPOT for an increasing number of users. Finally, we also investigate the throughput from the user perspective. The user throughput measures the rate at which tokens are received, as experienced by a single user. This metric is calculated by dividing the total number of response tokens by the duration from the request’s start ($t_{s}$) to the completion of the answer ($t_{e}$). We average this value across all users and report the mean and standard deviation.
@@ V-B System-Centric Metrics
Figure 4: Median Time to First Token: Baseline models (A,B,C) show lower TTFT values compared to 70B Expert Router model (D) due to the benefit of tensor-parallelism in the prefill phase. The 8B Expert Router configuration (E) shows similar TTFT to tensor-parallized models due to reduced computational demands. The whiskers on the right box plot show that the Expert Router models have higher minimum values, highlighting the extra latency from the routing gateway.
In addition to traditional metrics, we use a method for evaluating the system throughput over time by analyzing the recorded token timestamps. In each trial, we establish K time intervals and count the number of tokens present within each interval. Thus, we only count how many tokens the system produces in a specific time interval, regardless of to which user a specific token belongs. This strategy for measuring the throughput allows us to concentrate on the distribution of token timestamps, enabling us to analyze the system throughput dynamics. To make the different systems more easily comparable we calculate the mean throughput across all time intervals and present these values for specific levels of user concurrency in Section VI-B.
@@ VI Results
In this section, we present the outcomes of our experiments, analyzing the characteristics of the Expert Router and baseline configurations from both user and system perspectives. We systematically increased the number of concurrent users sending requests to the system to assess the performance under varied load conditions.
In addition to the performance metrics presented in Section V, we have also measured the latency introduced by the routing gateway. The average latency for each user when sending prompts via the routing gateway is 442 ms with a standard deviation of 302 ms. This latency remains stable irrespective of the number of concurrent users. The high standard deviation is caused by outliers with very high latencies, which could originate from a bottleneck on the DGX system during peak load times. The latency of the routing gateway will be considered in the further discussion of the results. As presented in Section IV-A, we will provide results for the Expert Router configurations under two different load scenarios: uniform and normal distributions. We will directly compare the two workload scenarios for the user throughput and the p99 response time. For all other metrics, we will focus on the results from the normal distribution, serving as an upper bound for latencies and a lower bound for throughput.
@@ VI-A User-Centric Metrics
Figure 5: Median Time per Output Token: The 8B Expert Router configuration (E) maintains the lowest TPOT due to its smaller number of parameters. Baseline model C shows higher TPOT due to increased communication overhead with tensor parallelism.
Figure 6: User Throughput and p99 Response time: Expert Router models (D, E) show a minimal difference between normal (Nor.) and uniform (Un.) workload distributions in both metrics, demonstrating robustness to varying workloads. In both metrics, model C and Expert Router configuration E show a clear advantage.
Figure 7: System throughput: Expert Router configuration E matches the throughput of baseline model C at 200 and 600 concurrent users, but falls behind at 1,000 due to larger TTFT values. Configuration D consistently shows the lowest peak and average throughput, highlighting the trade-off of using high-parameter configurations. The colored arrows indicate average throughput. The window size in which the tokens are counted is set to 2s.
Figure 4 illustrates the Time to First Token (TTFT) for an increasing number of users. As visible in the box plot on the right, the baseline models (A, B, C) consistently achieve lower minimum TTFT values than the Expert Router models (D, E). This is caused by the additional latency introduced by the routing gateway in the Expert Router architecture. However, this added latency does not result in a significant bottleneck during inference, as visible by the constant relative difference in the median TTFT values in the left plot. Baseline model C, which uses two tensor-parallelized 70B models, consistently exhibits the lowest TTFT throughout all levels of concurrency.
The better performance of the tensor-parallelized baseline models (A, B, C) can be attributed to their ability to efficiently distribute computations during the prefill phase of inference across multiple GPUs. In this phase, the input’s K and V vectors are computed and stored in the KV-cache. The parallel execution of the underlying matrix-matrix multiplications allows the tensor-parallelized baseline models to produce the first output token faster, leading to a lower overall TTFT [27]. The 8B model requires fewer computations, leading to similar TTFT values for configuration E as the tensor-parallelized models.
The advantage of less computation for the smaller 8B Expert Router configuration (E) becomes even more visible in Figure 5. Configuration E shows a clear advantage in the Time per Output Token (TPOT) compared to all other models. This is further highlighted by the small variance of the TPOT for high levels of user concurrency, visible in the box plot on the right in Figure 5. The elevated TPOT for the baseline models is due to the increased communication overhead required for the tensor-parallelization.
The results suggest that for moderate levels of user concurrency, the number of parameters of the models in the Expert Router configuration does not drastically affect the TTFT. Furthermore, even with many concurrent users, the TPOT remains within acceptable ranges for both Expert Router configurations. Despite the differences in TTFT, these findings demonstrate that the Expert Router configurations can deliver reliable performance across varied workloads.
Figure 6 I) evaluates the user throughput in response to increasing system load. This plot includes two result curves for each Expert Router configuration, representing requests distributed by normal (Nor.) and uniform (Un.) distributions across the clusters. For both Expert Router configurations, the difference between the two workload scenarios is minimal. This suggests that Expert Router configurations do not require perfectly balanced model utilization, and peaks in user requests have a limited effect on the user-perceived performance. Expert Router model E shows the highest user throughput up to a concurrency level of 800 users, highlighting its competitive performance from a user perspective.